OpenStudy (anonymous):

solve the linear equation. where necessary , choose no solution or all real number

OpenStudy (anonymous):

@Abbles

OpenStudy (anonymous):

Whats the equation?

OpenStudy (anonymous):

2(h+8)-h=h+16

OpenStudy (anonymous):

Alright first, distribute the 2

OpenStudy (anonymous):

2h+16-h+h+16

OpenStudy (abbles):

Distribute first. 2(h+8)-h=h+16 2h + 16 - h = h + 16 combine like terms h + 16 = h + 16 isolate the variable 0 = 0 The answer is all real numbers I believe. Because you can plug any number into the equation and it will solve.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Yea 0 = 0 means all real numbers